To determine the concentration of a solution of nitric acid, a 150.0-mL sample is placed in a flask and titrated with a 0.1175 M solution of barium hydroxide. A volume of 31.46 mL is required to reach the phenolphthalein endpoint. Calculate the concentration of nitric acid in the original sample.


_______ M.
1
Expert's answer
2019-03-14T04:00:59-0400

x mol 0.0037 mol

2HNO3 + Ba(OH)2 = Ba(NO3)2 + 2H2O

2 mol 1 mol


n(Ba(OH)2) = C(Ba(OH)2) * V(Ba(OH)2) = 0.1175 M * 0.03146 L = 0.0037 mol.

n(HNO3) = x = 0.0037 *2 / 1 mol = 0.00185 mol.

C(HNO3) = n(HNO3) / V(HNO3) = 0.00185 mol / 0.15 L = 0.0123 M.